OpenStudy (anonymous):

function of function 3 is 2and fuction of fuction2 is 1 wat will be the value of fuction0

OpenStudy (anonymous):

f(f(3))=2 f(f(2)=1 f(0)

OpenStudy (allank):

Hint: The function f(x) seems to be f(x)=x-0.5

OpenStudy (anonymous):

how do u know?

OpenStudy (anonymous):

i have the ans but i dont know how to solve

OpenStudy (anonymous):

if u can guide me throug the steps that would be great

OpenStudy (allank):

Sure. Look at the values: f(f(3)) would mean f(3)=3-0.5=2.5, then plug that 2.5 into the function again to get f(2.5)=2. Doing the same with the second case works, so the f(x)=x-0.5

OpenStudy (anonymous):

where did u get -0.5

OpenStudy (allank):

It's more of trial and error. Just thought about it and tested some possibilities.

OpenStudy (anonymous):

oh is that the only way?

OpenStudy (allank):

In this case, I believe so :)

OpenStudy (anonymous):

thanks :)

OpenStudy (allank):

Sure thing. You're welcome.